# A Constraints-Based Approach to Fully Interpretable Neural Networks for Detecting Learner Behaviors

### v1

@@ 1 Introduction
We specifically focus on the detection of gaming-the-system behavior, a type of learner disengagement with an educational task. We chose this due to the existence of previous expert-based features and models [17] that serve as a baseline of both accuracy and interpretability.
@@ 2.3 Interpretable gaming behavior detection
Through cognitive task analysis—in which an expert in the behavior of interest reasons through various action sequences and explains their thinking out loud—[17] were able to identify a set of features designed to capture elements of students’ problem-solving behavior that experts pay attention to when looking for gaming-the-system behavior. For example, these features included whether a student reuses the same answer in a different part of the problem interface or if the student enters consecutive similar answers. Each feature is binary, indicating whether it was observed or not for a given action.
@@
An example of one of these patterns can be seen in Figure 1. Each row is a different feature identified via cognitive task analysis, and each column indicates a student action (i.e. a different point in time). The pattern is a 3-action sequence. As originally written out, this pattern is as follows:
@@
. Refer to caption
@@ 3 Methods
For the sake of simplicity, we only use student action sequences from the dataset that have a fixed length of 5 actions (85% of all sequences). We use the same held-out test set as [17], consisting of 25% of the data. We further split the remaining data into a training/validation set using an 80/20 stratified split, using the validation set to tune hyperparameters and further refine our learned patterns. The final sets consist of 5,249 training clips, 1,313 validation clips, and 2,170 test clips, each with about 6% positively labeled instances.
Since convolutional filters must have a fixed kernel length (number of actions), we use a kernel length of 3, corresponding with the most common pattern length identified by the human expert. Theoretically, the model could learn patterns with smaller lengths than this, but not longer. We set the convolutional layer’s padding to 1 to allow the model to learn shorter patterns on the edges of input action sequences, and we set the number of filters to 2,048 to allow the model to learn a wide variety of patterns.
@@ 3.1.1 Binarize convolutional filter weights
The first regularization term constrains the model’s convolutional layer to learn weights very close to 00 or $1$ by penalizing values that deviate from these. Constraining the weights in this way enables a more straight-forward interpretation of each filter as a sequential pattern that emulates the binary presence or absence of specific student actions. This binarizing effect can be seen in the difference between the two left-most filters in Figure 3.
@@
where $W\in\mathbb{R}^{M\times k\times d}$ is the weight tensor for the convolution layer, $M$ is the number of filters (i.e. the number of patterns to learn), $k$ is the number of actions (sequence steps), and $d$ is the number of features.
@@
. Refer to caption
@@
3.1.2 Minimize positive weights per action
An additional regularization term places a penalty on actions that have a high number of positive weights. We introduced this term to encourage the model to focus only on the most relevant features per action step. We reasoned that having fewer activated features would also improve intelligibility by reducing cognitive load at explanation time.
@@ 3.1.3 Force single submission type
Note that this is slightly simplified from the original features described in [17], which used two separate incorrect attempt labels: wrong and bug. Some of the expert patterns also use attempt to indicate any submission type that is not help. We removed this feature to simplify the submission types—all expert patterns that used it also included a non-help-related feature in the same action, making it unnecessary.
@@
Figure 3 again shows the impact of this regularization term. The fourth filter from the left (where this constraint has been introduced) contains at most one positive submission type per action (column), unlike those to its left which don’t have this constraint.
@@ 3.1.4 Ensure possible feature combinations
Our final regularization term encourages the model to learn only positive weights in one of two feature sets for any single action. For example, the nature of our feature set makes it so that the first five features are help-related features, so they can only be positive when the submission type is help (third feature from the bottom in the filters visualized in Figure 3). By extension, all other features (except submission types) are attempt-related features and can only be activated when the submission type is NOT help.
This means that the input data will never have a positive feature in both feature sets in the same student action. Filters that don’t follow this guideline will never lead to a perfect match, so our constraint is designed to avoid such ineffective patterns. Compare the right-most filter in Figure 3 with the others to see the effect of this constraint.
@@ 3.3 Progressive constraining
We found that when the binarizing constraint was introduced too early, the weights had difficulty changing, even when the loss was high due to the penalties from other constraints. We also found that when the explicit filter matching constraint (via the differentiable thresholding fully connected layer) was introduced too late, the model often struggled to later learn patterns that matched the inputs exactly. Ultimately, we found that the best results were achieved when we introduced the constraints in the following order: explicit filter matching, possible feature combinations, single submission type, minimum positive weights per action, and binary filters.
@@ 3.4 Training across multiple eras
We trained our model for a total of 50 eras, each consisting of 200 epochs. At the end of each era, we reinitialized the weights for empty filters (those with all 00 weights) and those with precision < 0.3. We also reset all $\gamma_{i}$ values, restarting their staggered warm-up periods. This approach allowed the model to learn entirely new gaming-the-system patterns or refine existing ones by once again allowing gradients to flow and giving the chance for the weights to escape local minima, all while still adhering to the constraints we set out to enforce. We did not reset the $\alpha$ scaling weight. Figure 4 shows the model’s progressive constraining across two eras.
@@
. Refer to caption
@@
One advantage of our minimalist architecture—one with learnable weights only in the convolutional layer—is that we could save these weights at the end of each era, retaining the model’s entire set of learned parameters for future use. This provided us a far larger set of gaming-the-system patterns than what the model could learn in a single era. By analyzing these patterns, we were able to identify the most effective ones, remove duplicates, and insert the most effective ones back into the model for final inference. We further describe this process and report our results in the next section.
@@ 4 Results
Our approach generated a total of 25,981 filters with precision above 0.3. However, only 1,359 of these were unique. Furthermore, because of the nature of convolutional filters, if the positive weights in one filter are also positive in another, even if the latter has more positive weights, the former will match on the same inputs as the latter. In this scenario, the first filter has captured a more general gaming-the-system pattern than the second.
@@
. Refer to caption
@@
Table 1: Performance metrics on various datasets.
4.0.1 Comparison with expert patterns
@@
We made this comparison using the Levenshtein edit distance, which measures the minimum number of single-value edits (in our case, substitutions from 00 to $1$ or vice versa) required to change one pattern into another. Because some expert patterns contained 2 or 4 actions (as opposed to our filters’ 3), we expanded them to all possible 3-action sequences for this analysis. From 2-action patterns we created two separate patterns: one with a blank action at the beginning and another with a blank action at the end. From 4-action patterns we also created two separate patterns: one encompassing the first three actions and another with the last three.
We found that the average distance between our filters and the expert patterns was 13.5, with a standard deviation of 2.6. This suggests that our model’s learned filters differed quite significantly from those identified by human experts. Our model did not learn any of the expert patterns on its own. The most similar patterns had an edit distance of 3 and 5, but only three pairs were this similar (shown in Figure 6).
@@
. Refer to caption
@@
One noticeable difference is that our model’s learned filters tended to have many more positive features than the expert patterns. The model’s filters had a mean of 12.0 positive features, with a range of 4 to 18 and a standard deviation of 2.3. This is in contrast to the expert patterns, which—when taking into account only those with three actions—had a mean of 4.9 positive features, with a range of 4 to 7 and a standard deviation of 1.1. The entire set of expert patterns had a mean of 5.1 positive features, with a range of 4 to 7 and a standard deviation of 1.0.
@@ 5.1 Explainability
. Refer to caption
@@ 5.2 Limitations and future work
Another, less insurmountable, limitation of the current model is that the convolutional filters only allow for a fixed number of actions. That is, the input may be of any length, but the patterns can only have at most three student actions. This is an inherent limitation of the architecture, since the tensor containing the convolutional weights forces a fixed kernel size. One possible solution to explore in future work would be to provide separate but parallel convolutional layers, one for each desired sequence length. The feature maps, or output of each layer, could then be concatenated and passed through the thresholding layer as normal.

### v2

@@ 1 Introduction
We specifically focus on the detection of gaming-the-system behavior, a type of learner disengagement with an educational task. We chose this due to the existence of previous expert-based features and models [17] that serve as a baseline of both accuracy and interpretability. We also chose to use a deep learning model due to their rising popularity among EDM researchers and the difficulty of faithfully interpreting their decision-making process [19].
@@ 2.3 Interpretable gaming behavior detection
Through cognitive task analysis—in which an expert in the behavior of interest reasons through various action sequences and explains their thinking out loud—[17] were able to identify a set of features designed to capture elements of students’ problem-solving behavior that experts pay attention to when looking for gaming-the-system behavior. For example, these features included whether a student reuses the same answer in a different part of the problem interface or if the student enters consecutive similar answers. Each feature is binary, indicating whether it was observed or not for a given action step. An action step is a point in time at which data was collected, and is triggered by the student either making a submission or asking for a hint.
@@
An example of one of these patterns can be seen in Figure 1. Each row is a different feature identified via cognitive task analysis, and each column indicates a student action step. The pattern is a 3-action sequence. As originally written out, this pattern is as follows:
@@
Example expert pattern from cognitive model. Refer to caption
@@
[…]
Model architecture. Refer to caption
Figure 2: Model architecture.
@@ 3 Methods
For the sake of simplicity, we only use student action sequences from the dataset that have a fixed length of 5 action steps (85% of all sequences). We use the same held-out test set as [17], consisting of 25% of the data. We further split the remaining data into a training/validation set using an 80/20 stratified split, using the validation set to tune hyperparameters and further refine our learned patterns. The final sets consist of 5,249 training clips, 1,313 validation clips, and 2,170 test clips, each with about 6% positively labeled instances.
Since convolutional filters must have a fixed kernel length (i.e. number of action steps), we use a kernel length of 3, corresponding with the most common pattern length identified by the human expert. Theoretically, the model could learn patterns with smaller lengths than this, but not longer. We set the convolutional layer’s padding to 1 to allow the model to learn shorter patterns on the edges of input action sequences, and we set the number of filters to 2,048 to allow the model to learn a wide variety of patterns.
@@ 3.1.1 Binarize convolutional filter weights
The first regularization term constrains the model’s convolutional layer to learn weights very close to 00 or $1$ by penalizing values that deviate from these. Constraining the weights in this way enables a more straight-forward interpretation of each filter as a sequential pattern that emulates the binary presence or absence of specific features per action step. This binarizing effect can be seen in the difference between the two left-most filters in Figure 3.
@@
where $W\in\mathbb{R}^{M\times k\times d}$ is the weight tensor for the convolution layer, $M$ is the number of filters (i.e. the number of patterns to learn), $k$ is the number of action steps, and $d$ is the number of features.
@@
Visualization of the effects of increasing constraints on the model’s convolutional filters. Refer to caption
@@
3.1.2 Minimize positive weights per action step
An additional regularization term places a penalty on action steps that have a high number of positive weights. We introduced this term to encourage the model to focus only on the most relevant features per action step. We reasoned that having fewer activated features would also improve intelligibility by reducing cognitive load at explanation time.
@@ 3.1.3 Force single submission type
Note that this is slightly simplified from the original features described in [17], which used two separate incorrect attempt labels: wrong and bug. Some of the expert patterns also use attempt to indicate any submission type that is not help. We removed this feature to simplify the submission types—all expert patterns that used it also included a non-help-related feature in the same action step, making it unnecessary.
@@
Figure 3 again shows the impact of this regularization term. The fourth filter from the left (where this constraint has been introduced) contains at most one positive submission type per action step (column), unlike those to its left which don’t have this constraint.
@@ 3.1.4 Ensure possible feature combinations
Our final regularization term encourages the model to learn only positive weights in one of two feature sets for any single action step. For example, the nature of our feature set makes it so that the first five features are help-related features, so they can only be positive when the submission type is help (third feature from the bottom in the filters visualized in Figure 3). By extension, all other features (except submission types) are attempt-related features and can only be activated when the submission type is NOT help.
This means that the input data will never have a positive feature in both feature sets in the same action step. Filters that don’t follow this guideline will never lead to a perfect match, so our constraint is designed to avoid such ineffective patterns. Compare the right-most filter in Figure 3 with the others to see the effect of this constraint.
@@ 3.3 Progressive constraining
We found that when the binarizing constraint was introduced too early, the weights had difficulty changing, even when the loss was high due to the penalties from other constraints. We also found that when the explicit filter matching constraint (via the differentiable thresholding fully connected layer) was introduced too late, the model often struggled to later learn patterns that matched the inputs exactly. Ultimately, we found that the best results were achieved when we introduced the constraints in the following order: explicit filter matching, possible feature combinations, single submission type, minimum positive weights per action step, and binary filters.
@@ 3.4 Training across multiple eras
We trained our model for a total of 50 eras, each consisting of 200 epochs. At the end of each era, we reinitialized the weights for empty filters (those with all 00 weights) and for those with precision < 0.3. We also reset all $\gamma_{i}$ values, restarting their staggered warm-up periods. This approach allowed the model to learn entirely new gaming-the-system patterns or refine existing ones by once again allowing gradients to flow and giving the chance for the weights to escape local minima, all while still adhering to the constraints we set out to enforce. We did not reset the $\alpha$ scaling weight. Figure 4 shows the model’s progressive constraining across two eras.
One advantage of our minimalist architecture—one with learnable weights only in the convolutional layer—is that we could save these weights to an external file at the end of each era, retaining the model’s entire set of learned parameters for future use. This provided us a far larger set of gaming-the-system patterns than what the model could learn in a single era. By analyzing these patterns, we were able to identify the most effective ones, remove duplicates, and insert those remaining back into the model for final inference. We further describe this process and report our results in the next section.
@@
Scaling weight increases with staggered warm-up periods across two eras of 200 epochs each. Refer to caption
@@ 4 Results
Our approach generated a total of 102,400 filters (2,048 per era for 50 eras) from which we saved the 25,981 filters that achieved a precision above 0.3. Because the non-reinitialized filters were carried over to the initial state in subsequent eras, there was a large amount of repetition in the total set of filters identified during training. After removing duplicate filters, only 1,359 were unique. Furthermore, because of the nature of convolutional filters, if the positive weights in one filter are also positive in another, even if the latter has more positive weights, the former will match on the same inputs as the latter. In this scenario, the first filter has captured a more general gaming-the-system pattern than the second.
@@
Cumulative metrics on patterns sorted descending by precision on the training set. Refer to caption
@@
4.1 Comparison with expert patterns
@@
Furthermore, because some expert patterns contained 2 or 4 action steps (as opposed to our filters’ 3), we expanded them to all possible 3-action sequences for this analysis. From 2-action patterns we created two separate patterns: one with a blank action step at the beginning and another with a blank action step at the end. From 4-action patterns we also created two separate patterns: one encompassing the first three action steps and another the last three. While this latter scenario likely leads to trimmed patterns that the expert may no longer consider indicative of gaming-the-system behavior, we reasoned that such a comparison with our model’s learned filters would still be informative.
We made our comparison using the Levenshtein edit distance, which measures the minimum number of single-value edits (in our case, substitutions from 00 to $1$ or vice versa) required to change one pattern into another. We found that the average distance between our filters and the expert patterns was 13.5, with a standard deviation of 2.6. This suggests that our model’s learned filters differed quite significantly from those identified by human experts. Our model did not learn any of the expert patterns on its own. The most similar patterns had an edit distance of 3 and 5, but only three pairs were this similar (shown in Figure 6).
@@
Expert patterns (red) and our model’s learned filters (green) for the most similar pairs. Refer to caption
@@
One noticeable difference is that our model’s learned filters tended to have many more positive features than the expert patterns. The model’s filters had a mean of 12.0 positive features, with a range of 4 to 18 and a standard deviation of 2.3. This is in contrast to the expert patterns, which—when taking into account only those with three action steps—had a mean of 4.9 positive features, with a range of 4 to 7 and a standard deviation of 1.1. The entire set of expert patterns had a mean of 5.1 positive features, with a range of 4 to 7 and a standard deviation of 1.0.
@@ 5.1 Explainability
Example explanations created from the model’s learned filters. Refer to caption
@@ 5.2 Limitations and future work
Another, less insurmountable, limitation of the current model is that the convolutional filters only allow for a fixed number of action steps. That is, the input may be of any length, but the patterns can only have at most three student action steps. This is an inherent limitation of the architecture, since the tensor containing the convolutional weights forces a fixed kernel size. One possible solution to explore in future work would be to provide separate but parallel convolutional layers, one for each desired sequence length. The feature maps, or output of each layer, could then be concatenated and passed through the thresholding layer as normal.
@@ 6 Conclusion
7 Acknowledgments
This study is supported by the National Science Foundation under Award #1942962. Any conclusions expressed in this material do not necessarily reflect the views of the NSF.